OpenStudy (anonymous):

How would you integrate\[\int\limits_{0}^{\infty}\Phi \sin \phi \Phi''d \phi\]

OpenStudy (anonymous):

what is \[\Phi \]

OpenStudy (anonymous):

Sorry,\[\Phi(\phi)\] Uppercase phi is a function of lowercase phi

OpenStudy (anonymous):

@mahmit2012

OpenStudy (anonymous):

can you state the original problem ?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Given\[\Phi''+\frac{\cos \phi}{\sin \phi}\Phi'+(\nu-\frac{n^2}{\sin^2 \phi})\Phi=0\]Notice that this is a bessel function of order v-n^2/sin^2(phi) Show that\[\nu>0\]We have to start by multiplying both sides by\[\Phi \sin \phi\]And integrate from 0 -> infinity

OpenStudy (anonymous):

I meant of order n^2 >.<

OpenStudy (anonymous):

lol...i got it...

OpenStudy (anonymous):

i think we can rearrange to the standard form with a variable change or something like this :\[t=\sqrt{\nu} \sin \phi\]

OpenStudy (anonymous):

Hm might play around with that, the professor literally just multiplied everything by\[\Phi \sin \phi\] and integrated.